TUNABLE VACUUM ULTRAVIOLET PHOTOIONIZATION MEASUREMENTS OF CARBON CLUSTERS

\underline{B. W. TICKNOR}, S. E. WHEELER, H. F. SCHAEFER III and M. A. DUNCAN, Department of Chemistry, Univeristy of Georgia, Athens, GA, 30602-2556; L. BELAU and M. AHMED, Chemical Sciences Division, Ernest Orlando Lawerence Berkeley National Laboratory, Berkeley, California 94720; S. R. LEONE, Departments of Chemistry and Physics and Ernest Orlando Lawrence Berkeley National Laboratory, Univerisity of California, Berkeley, California 94720.

The photoionization thresholds for neutral carbon clusters have been measured using tunable vacuum ultraviolet light at the Advanced Light Source (ALS). A laser ablation cluster source with a high repetition rate vaporization laser is coupled to the quasi-contionous light from the ALS. Ionized clusters are detected as a function of photon wavelength to determine the onset of ionization. Ionization thresholds for carbon clusters up to C_{15} have been determined. These thresholds are compared to ionization potentials calculated from ab initio methods (extroplated CCSD(T) energies computed at cc-pVTZ ROCCSD(T) optimized geometries), which allows the investigation of isomeric structures in the molecular beam for some cluster sizes.